# 0pt[Uncaptioned image] Dual loops     0pt[Uncaptioned image] Polycube     0pt[Uncaptioned image] Polycube segmentation

### v3

@@ Abstract
Polycube segmentations for 3D models effectively support a wide variety of applications such as seamless texture mapping, spline fitting, structured multi-block grid generation, and hexahedral mesh construction. However, the automated construction of valid polycube segmentations suffers from robustness issues: state-of-the-art methods are not guaranteed to find a valid solution. In this paper we present an iterative algorithm which is guaranteed to return a valid polycube segmentation for 3D models of any genus. Our algorithm is based on a dual representation of polycubes. Starting from an initial simple polycube of the correct genus, together with the corresponding dual loop structure and polycube segmentation, we iteratively refine the polycube, loop structure, and segmentation, while maintaining the correctness of the solution. Our algorithm is robust by construction: at any point during the iterative process the current segmentation is valid. Furthermore, the iterative nature of our algorithm facilitates a seamless trade-off between quality and complexity of the solution. Our algorithm can be implemented using comparatively simple algorithmic building blocks; our experimental evaluation establishes that the quality of our polycube segmentations is on par with, or exceeding, the state-of-the-art.
@@ 1 Introduction
Polycubes are orthogonal polyhedra with axis-aligned quadrilateral faces. The simple structure of polycubes enables efficient solutions to various challenging geometric problems. Bijective mappings from general shapes to polycubes, known as polycube maps, enable the transfer of solutions computed on polycubes to those general shapes. Polycube maps are used to solve problems such as texture mapping [THCM04], spline fitting [WHL07], multi-block grid generation [BBG09], and hexahedral meshing [PCS22].
@@
Formally, a polycube map $f$ is a continuous map from a polycube $Q$ of genus $g$ to a closed $2$-dimensional surface $\mathcal{M}$ of the same genus. The edges of $Q$ map to a segmentation of $\mathcal{M}$ into patches that correspond to the faces of $Q$, known as a polycube segmentation (see Figure Robust Construction of Polycube Segmentations via Dual Loops). Each patch has a label ±$\pm X/\pm Y/\pm Z$: the direction of the normal vector of its corresponding polycube face.
@@
Since their introduction in 2004 [THCM04], numerous methods have been proposed for constructing polycube maps. The quality of a polycube map is determined by two competing factors: the complexity of the polycube and the distortion of the mapping. Methods that construct polycube maps must find a balance between these conflicting factors. The state-of-the-art achieves a good trade-off between low distortion and low complexity [GSZ11, LVS13, DPM22].
@@
A common approach for constructing polycube maps is to directly construct a polycube segmentation [DPM22, FBL16, GSZ11, HZ16, HZL17, LVS13, ZLL17]. We call this the primal approach. Developing a polycube segmentation into a full (surface or volume) polycube map can be done via various established methods [GSZ11, PRR22].
The proposed methods for constructing polycube segmentations  [SR15, ZLW19, PRR22, MPBL23] suffer from various robustness issues. The methods incorrectly identify invalid segmentations as valid, and vice-versa, since they can not guarantee that the constructed polycube segmentations are valid. The existing characterizations of polycubes [EM10, HLW24, ZLW19], used for checking whether a segmentation corresponds to a valid polycube, are neither necessary nor sufficient [MPBL23, Sok16, SSV25]. As a result, current primal methods unnecessarily restrict their solution space or generate invalid solutions.
@@
In this paper we describe an iterative robust algorithm to construct polycube segmentations. Our algorithm is based on a recent characterization of polycubes via its dual loop structure [SSV25]. This characterization is complete, but characterizes a specific subset of polycubes: those with strictly quadrilateral faces. This does not necessarily restrict the space of shapes we can form, but requires us to construct such polycube segmentations in order to use their characterization. The characterization is not directly applicable to the previously mentioned approaches, as they construct polycube segmentations with arbitrarily complex patches. Note that we can turn quad-only polycube segmentations into the arbitrarily-complex polycube segmentations by simply "hiding" the so-called flat edges, edges between two patches with the same label.
@@
We start with a simple loop structure that corresponds to a polycube that has the same genus as the input surface $\mathcal{M}$. Then we iteratively add or remove loops to refine or simplify the loop structure in a way that preserves validity. The iterative nature of our algorithm allows us to explicitly control the trade-off between the complexity of the polycube and the distortion of the mapping. We review related work in Section 2 and give the necessary mathematical background, including the dual representation of polycubes, in Section 3. In Section 4 we describe our algorithm in detail. In Section 5 we evaluate our algorithm experimentally. Finally, we conclude with avenues for future work in Section 6.
@@
Figure 1: Our iterative process, akin to the development of an egg to a guppy to an adult fish.
@@ 2 Related Work
The most prominent methods for polycube map construction are based on the deformation approach by [GSZ11] which was later improved or extended by many [LVS13, HJS14, YZWL14, YFL19, FBL16, DPM22, ZLW19, FXBH16, GLYL20, MCBC22, CLS16]. The deformation-based approaches find an initial polycube segmentation in various ways, none of which guarantees that the segmentation corresponds to a valid polycube. Then, the input shape is gradually deformed until the surface faces are oriented toward their assigned target label. The initial polycube labeling may be adjusted during the deformation process. Many deformation methods use the characterization of orthogonal polyhedra by [EM10] to increase their robustness. However, this characterization is neither sufficient nor necessary for valid polycubes [SR15, MCBC22, PCS22, MPBL23].
An alternative set of methods is based on voxelization [WYZ11, YFL19, YZWL14]. In general, a polycube map can be obtained by projecting a voxelized version of $\mathcal{M}$ back onto $\mathcal{M}$. Although voxelizing $\mathcal{M}$ yields a valid polycube, the projection cannot be done robustly and may fail to preserve the topology of $\mathcal{M}$. Furtghermore, voxelization results in overly detailed polycubes, which is not desirable for most downstream applications.
Finally, since polycube maps are popular for solving the hexahedral meshing problem, some methods focus on optimizing specifically for hexahedral meshing. These methods obtain more manageble input surfaces that result in higher quality hexahedral meshes by introducing cuts or deforming the input surface [FXBH16, GLYL20, MCBC22]. While this works for hexahedral meshing of the interior volume of shapes, the methods are not suitable for other downstream applications (such as multi-block grid generation) where preservation of topology and low distortion with the original input model are crucial.
A different line of work by Baumeister and Kobbelt [BK23] and Snoep et al.[SSV25] (building on insights from Biedl and Genc[BG04]), takes a dual approach. Rather than directly constructing polycube segmentations, they first establish a valid dual structure on $\mathcal{M}$ that is then primalized into a guaranteed valid polycube segmentation. This idea mirrors the dual approach in quad meshing [CBK12]. While experiments suggest this method can work, the practical aspects of implementing a complete pipeline, along with its limitations and how it compares to existing solutions, have yet to be fully explored.
@@ 3 Polycubes via Dual Loops
The algorithm in this paper is based on the characterization of polycubes by Snoep, Speckmann, and Verbeek [SSV25]. In this section we review the necessary definitions and results from their work.
@@ 3.1 Dual structure
If you consider the normals of the quadrilateral faces of a single loop, they always correspond to exactly two axes ($\{X,Y,Z\}$). As such, the loops can labeled as either an $X$-, $Y$-, or $Z$-loop as follows: an $X$-loop traverses faces with normals aligned to $Y$- and $Z$-axes. The $Y$- and $Z$-loops are similarly defined.
@@
We can use the full set of $X$-loops of an oriented loop structure to partition the underlying space (surface or polycube) into regions. We refer to these regions as $X$-zones. We can use these zones to capture the global structure of the polycube (with regard to the X-axis) in a directed graph. Specifically, the $X$-graph has a vertex for each $X$-zone, and a directed edge $(u,v)$ for every $X$-loop with $u$ and $v$ corresponding to the $X$-zones on the negative and positive side of the loop, respectively. We can define the $Y$-zones, $Z$-zones, $Y$-graph and $Z$-graph similarly.
@@ 3.2 Characterization
Consider a loop structure embedded on some arbitrary surface (not necessarily a polycube). In the remainder of this paper we implicitly assume that each loop in a loop structure is both oriented (defining negative and positive sides) and labeled with either $X$, $Y$, or $Z$, unless stated otherwise. There exists a complete set of rules that characterize which of these loop structures correspond to polycube loop structures; that is, the loop structures that are dual to a polycube.
@@
    The $X$-, $Y$-, and $Z$-graphs are acyclic.
The following result was shown by Snoep et al.
Theorem 1 ([SSV25])
@@ 4 Algorithm
In this section we present our algorithm for computing polycube segmentations of triangulated 3D models. We first give a general overview of the algorithm, before we dive deeper into the most important parts of the algorithm: adding loops to a polycube loop structure, computing the polycube segmentation from a polycube loop structure, and optimizing the polycube loop structure.
The input to our algorithm is a 3D model represented as a triangulated surface mesh $\mathcal{M}=(V,T)$, where $V$ is the set of vertices and $T$ is the set of triangular faces, each defined by a triplet $(v_{i},v_{j},v_{k})$ of vertices. The edges of $\mathcal{M}$ are implicitly defined by the faces. We assume that $\mathcal{M}$ is a well-formed, orientable, manifold surface of arbitrary genus that bounds a single volume. Additionally, $\mathcal{M}$ is embedded in $\mathbb{R}^{3}$, meaning that every vertex $v\in V$ has an associated position $p(v)=(x(v),y(v),z(v))\in\mathbb{R}^{3}$, and each triangle $t\in T$ has a corresponding normal vector $n(t)$.
@@
Our algorithm does not compute a polycube segmentation directly, but instead focuses on computing a valid (polycube) loop structure $\mathcal{L}$. The loops of $\mathcal{L}$ are directly embedded on the surface mesh $\mathcal{M}$. By ensuring that $\mathcal{L}$ satisfies the properties of Definition 3, we can guarantee that we can compute a polycube segmentation from $\mathcal{L}$ at any step of the algorithm. Our algorithm now proceeds as follows. First, we initialize a loop structure of minimum complexity on $\mathcal{M}$. For a model of genus zero, this is simply a loop structure consisting of three loops (one for each axis label), which corresponds to a single cube as polycube. For models of higher genus, this initialization is more involved (see Section 4.3 for more details). Next, we alter the loop structure by adding or removing loops. In Section 4.1 we explain in detail how to find good loops that improve the quality of the polycube segmentation, and at the same time preserve the characterization of a polycube loop structure stated in Definition 3. Once we have obtained a good loop structure, we compute a corresponding polycube segmentation by placing a corner in every loop region, and connecting the corners of adjacent loop regions via non-intersecting paths on $\mathcal{M}$. We refer to this process as primalization and discuss this step in more detail in Section 4.2. To obtain the best loop structure $\mathcal{L}$, we actually maintain a collection of loop structures, and we utilize a type of evolutionary algorithm to optimize the quality of the corresponding polycube segmentations (see Section 4.3 for details).
Observe that our algorithm does not actually compute the corresponding polycube of the polycube segmentation. However, our approach makes it relatively straightforward to compute the corresponding unique polycube (up to order-equivalence), and hence we often include the corresponding polycubes in our figures for clarity. Since this is not the main purpose of our algorithm, we will not describe how to compute these polycubes, and we will not try to optimize their exact geometric representation.
@@
To find the best loop structure via our evolutionary algorithm, we need to define the quality of a polycube segmentation. There is no single standard for measuring the quality of a polycube segmentation, as different applications have different requirements. Polycube maps are commonly used in hexahedral meshing, spline fitting, and structured multi-block grid generation, each prioritizing different aspects such as minimal distortion or fewer patches in the segmentation. In general, existing approaches seek to balance complexity, often tied to the number of vertices in the corresponding polycube, and distortion, which captures how well the polycube map transforms the original surface into the polycube. In this paper we specifically choose to optimize the following two criteria.
Alignment:
@@
This aspect captures how well the segmentation aligns with the original mesh geometry. Specifically, we use the angular deviation between the normal of each triangle and its assigned label.
Complexity:
This aspect captures the complexity of the polycube segmentation. Since our algorithm focuses on computing a loop structure, we measure the complexity via the number of loops in the loop structure. Although there is no direct relation between the number of loops and the complexity of the resulting polycube segmentation, the number of loops acts as a good proxy.
In Section 4.3 we discuss which exact metrics we use and how we handle the resulting multi-objective optimization problem. Note that the alignment cannot directly be computed on a loop structure, but it depends on a polycube segmentation. Therefore, whenever our algorithm evaluates the quality of a loop structure, it has to compute a corresponding polycube segmentation through the primalization step.

### v4

@@ Abstract
Polycube segmentations for 3D models effectively support a wide variety of applications such as seamless texture mapping, spline fitting, structured multi-block grid generation, and hexahedral mesh construction. However, the automated construction of valid polycube segmentations suffers from robustness issues: state-of-the-art methods are not guaranteed to find a valid solution. In this paper we present DualCube: an iterative algorithm which is guaranteed to return a valid polycube segmentation for 3D models of any genus. Our algorithm is based on a dual representation of polycubes. Starting from an initial simple polycube of the correct genus, together with the corresponding dual loop structure and polycube segmentation, we iteratively refine the polycube, loop structure, and segmentation, while maintaining the correctness of the solution. DualCube is robust by construction: at any point during the iterative process the current segmentation is valid. Its iterative nature furthermore facilitates a seamless trade-off between quality and complexity of the solution. DualCube can be implemented using comparatively simple algorithmic building blocks; our experimental evaluation establishes that the quality of our polycube segmentations is on par with, or exceeding, the state-of-the-art.
volume: 44issue: 5
@@ 1 Introduction
Polycubes, or orthogonal polyhedra with axis-aligned quadrilateral faces, enable efficient approaches to various challenging geometric problems. Polycube maps are bijective mappings from general shapes to polycubes. They allow for the transfer of efficient solutions computed on polycubes to those general shapes. As such, polycube maps are used to solve texture mapping [THCM04], spline fitting [WHL07], multi-block grid generation [BBG09], and hexahedral meshing [PCS22].
@@
Formally, a polycube map $f$ is a continuous function from a polycube $Q$ of genus $g$ to a closed 2-dimensional surface $\mathcal{M}$ with the same genus. The edges of $Q$ mapped onto $\mathcal{M}$ define a segmentation of patches, where each patch corresponds to face of $Q$. This segmentation is called a polycube segmentation (see Figure Robust Construction of Polycube Segmentations via Dual Loops). Each patch has a corresponding label $\{+X,-X,+Y,-Y,+Z,-Z\}$, indicating the direction of the normal vector of the mapped face. The construction of a valid polycube segmentation is a common prerequisite for generating high-quality polycube maps. However, current methods fail to ensure the validity of the segmentation.
@@
Since their introduction in 2004 [THCM04], numerous methods have been proposed for constructing polycube maps. The quality of a polycube map is determined by two competing factors: the complexity of the polycube and the distortion of the mapping. Methods that construct polycube maps must balance these conflicting factors. The state-of-the-art achieves a good trade-off between low distortion and low complexity [GSZ11, LVS13, DPM22].
@@
A common approach for constructing polycube maps is to first construct a polycube segmentation. Developing a polycube segmentation into a full (surface or volume) polycube map can be done via various established methods [GSZ11, PRR22]. Most current approaches construct a polycube segmentation directly [DPM22, FBL16, GSZ11, HZ16, HZL17, LVS13, ZLL17], we call this the primal approach. These construction methods suffer from various robustness issues [MPBL23, PRR22, SR15, ZLW19]. The methods incorrectly identify invalid segmentations as valid, and vice-versa, since they can not guarantee that the constructed polycube segmentations are valid.
The methods rely on a characterization of orthogonal polyhedra [EM10], used for checking whether a segmentation corresponds to a valid polycube. This characterization [EM10], and its extensions [HLW24, ZLW19] are neither necessary nor sufficient [MPBL23, Sok16]. As a result, the primal methods unnecessarily restrict their solution space or generate invalid solutions. We recently introduced a complete characterization of polycubes of arbitrary genus via dual loops [SSV25] by extending work of Biedl and Genc [BG04] and Baumeister and Kobelt [BK23]. Currently, no algorithm exists that utilizes the dual characterization of polycubes and the accompanying loop structure to construct high-quality polycube segmentations.
@@
In this paper we describe DualCube: an iterative robust algorithm to construct polycube segmentations based on our characterization of polycubes via their dual loop structure [SSV25]. This characterization is complete, but characterizes a specific subset of polycubes: those with strictly quadrilateral faces. As such this characterization cannot be used directly with the primal methods, since they construct polycube segmentations with arbitrarily complex patches. Note that we can turn quad-only polycube segmentations into arbitrarily-complex polycube segmentations by simply ignoring the so-called flat edges, edges between two patches with the same label.
@@
The iterative nature of our algorithm allows us to explicitly control the trade-off between the complexity of the polycube and the distortion of the mapping, see Figure 1. First, we embed a simple loop structure that corresponds to a polycube that has the same genus as the input surface $\mathcal{M}$. Then we iteratively add or remove loops to refine or simplify the loop structure in a way that preserves validity. The embedded loop structure can be transformed into a polycube segmentation through a primalization step.
We review related work in Section 2. We discuss the necessary mathematical background, including the dual representation of polycubes, in Section 3. We describe our algorithm, including all technical details, in Section 4. We evaluate DualCube experimentally and showcase the results in Section 5. Finally, we conclude with avenues for future work in Section 6.
@@
Figure 1: Iterative construction of a polycube segmentation, akin to the development of an egg to a guppy to an adult fish.
@@ 2 Related Work
The most prominent methods for polycube map construction are based on the deformation approach by [GSZ11] which was later improved or extended by many [CLS16, DPM22, FXBH16, FBL16, GLYL20, HJS14, LVS13, MCBC22, YFL19, YZWL14, ZLW19]. The deformation-based approaches find an initial polycube segmentation in various ways, none of which guarantees that the segmentation corresponds to a valid polycube. Then, the input shape is gradually deformed until the surface faces are oriented toward their assigned target label. The initial polycube labeling may be adjusted during the deformation process. These methods use the characterization of orthogonal polyhedra by [EM10], which is neither sufficient nor necessary for valid polycubes [MPBL23, SR15].
Another set of methods is based on voxelization [WYZ11, YFL19, YZWL14]. In general, a polycube map can be obtained by projecting a voxelized version of $\mathcal{M}$ back onto $\mathcal{M}$. Although voxelizing $\mathcal{M}$ yields a valid polycube, the projection cannot be done robustly and may fail to preserve the topology. In addition, voxelization often leads to overly complex polycubes, which is not desirable for most downstream applications.
Finally, since polycube maps are popular for solving the hexahedral meshing problem, some methods focus on optimizing specifically for hexahedral meshing. These methods transform the input surfaces into more manageable shapes that result in higher quality hexahedral meshes by introducing cuts or deforming the input surface [FXBH16, GLYL20, MCBC22]. While these techniques are effective for interior volumetric meshing, such as hexahedral meshing, they may heavily distort the original surface or apply topological simplifications such as the cutting of handles. These modifications may preserve the volume but alter the original surface. In contrast, exterior volumetric meshing applications, such as multi-block grid generation, impose stricter constraints: the output must remain topologically and geometrically consistent with the input.
A different line of work started by Biedl and Genc [BG04, BG09] and Baumeister and Kobbelt [BK23] takes a dual approach, characterizing polycubes through their dual structures consisting of loops traversing the quadrilateral faces of a polycube. We recently introduced a complete characterization of polycubes of arbitrary genus via these dual loops [SSV25]. In the same paper, we briefly outlined a proof-of-concept algorithm that first constructs a valid dual structure, which is then primalized into a guaranteed valid polycube segmentation, inspired by the dual approach used for quad meshing [CBK12]. However, our prior work focused on theoretical foundations and did not include technical details, algorithmic analysis, and experimental evaluation against the state-of-the-art.
@@ 3 Polycubes via Dual Loops
Our algorithm builds on our characterization [SSV25] of polycubes via dual loops: loops that correspond to axis-aligned strips of quadrilateral faces. In the following we review the necessary mathematical background and the salient parts of our characterization.
@@ 3.1 Dual structure
If you consider the normals of the quadrilateral faces of a single loop, they always correspond to exactly two axes ($\{X,Y,Z\}$). As such, the loops can be labeled as either an $X$-, $Y$-, or $Z$-loop as follows: an $X$-loop traverses faces with normals aligned to $Y$- and $Z$-axes. The $Y$- and $Z$-loops are similarly defined.
@@
We can use the full set of $X$-loops of an oriented loop structure to partition the underlying space (surface or polycube) into regions. We refer to these regions as $X$-zones. We can use these zones to capture the global structure of the polycube (with regard to the X-axis) in a directed graph. Specifically, the $X$-graph has a vertex for each $X$-zone, and a directed edge $(u,v)$ for every $X$-loop with $u$ and $v$ corresponding to the $X$-zones on the negative and positive side of the loop, respectively. We can define zones and graph for $Y$ and $Z$ similarly. The $X$-, $Y$-, and $Z$-graph are also called the level graphs.
@@ 3.2 Characterization
Consider a loop structure embedded on some arbitrary surface (not necessarily a polycube). In the remainder of this paper we implicitly assume that each loop in a loop structure is both oriented (defining negative and positive sides) and labeled with either $X$, $Y$, or $Z$, unless stated otherwise. There exists a complete set of rules that characterize which of these loop structures correspond to polycube loop structures; those that are dual to a polycube.
@@
    The level graphs are acyclic.
We previously proved the following theorem:
Theorem 1 (Snoep, Speckmann, and Verbeek [SSV25])
@@
3.3 Valid modifications
Figure 4: Loop additions (left to right) in the dual of a polycube. Loop removals can be seen as the reverse of addition (right to left).
Our dual characterization of polycubes [SSV25] supports two basic valid operations that locally transform one valid structure into another: loop addition and loop removal. See Figure 4. These operations are sufficient for constructing many, but not all, polycube configurations. It remains unclear whether they suffice for reachability between all genus-zero polycube configurations [SSV25]. For higher-genus cases, counterexamples exist: for example, adding or removing a single loop cannot change the orientation of a hole, making configurations with differing hole orientations unreachable from one another using only (single) loop additions and removals.
Loop addition.  Loop addition involves inserting a new loop into an existing structure such that the modified configuration remains a valid polycube loop structure. To do this, the authors define three directed graphs that encode permissible transitions between loop segments based on local constraints of Definition 3.
Given a loop structure $\mathcal{L}$, a base graph $G_{L}$ is constructed with one vertex per loop segment and edges between segments that share a loop region. From this, three filtered graphs $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$ are derived, corresponding to candidate loops with labels $X$, $Y$, or $Z$, respectively. Each filtered graph retains only the edges whose insertion into the loop structure results in valid loop regions, according to Conditions 1–4 of Definition 3. A candidate loop is considered valid if and only if it forms a non-trivial, simple (visits a loop region at most once), directed cycle in its corresponding graph $G_{\Delta}^{V}$, where $\Delta\in\{X,Y,Z\}$. In [SSV25] we show that such loop additions do not introduce cycles in the level graphs, thus also preserving Condition 5.
Loop removal.  Loop removal is the inverse operation, in which an existing loop $\lambda$ is removed from the loop structure. Let $\lambda$ be an $X$-loop, its removal merges each pair of loop regions separated by loop $\lambda$. Conditions 1–4 of Definition 3 can be verified locally at all loop regions adjacent to $\lambda$. To ensure that Condition 5 is preserved, the resulting $X$-graph must be checked for cycles.
@@ 4 Algorithm
In this section we present DualCube: our algorithm for computing polycube segmentations of triangulated 3D models. We first give a general overview, before we dive deeper into the technical details for modifying the polycube loop structure, constructing the polycube segmentation from a polycube loop structure, and optimizing the polycube loop structure.
The input to DualCube is a 3D model represented as a triangulated surface mesh $\mathcal{M}=(V,T)$, where $V$ is the set of vertices and $T$ is the set of triangular faces, each defined by a triplet $(v_{i},v_{j},v_{k})$ of vertices. The edges of $\mathcal{M}$ are implicitly defined by the faces. We assume that $\mathcal{M}$ is a well-formed, orientable, manifold surface of arbitrary genus that bounds a single volume. Additionally, $\mathcal{M}$ is embedded in $\mathbb{R}^{3}$, meaning that every vertex $v\in V$ has an associated position $p(v)=(x(v),y(v),z(v))\in\mathbb{R}^{3}$, and each triangle $t\in T$ has a corresponding normal vector $n(t)$.
@@
Our algorithm does not compute a polycube segmentation directly, but instead focuses on computing a valid (polycube) loop structure $\mathcal{L}$. The loops of $\mathcal{L}$ are directly embedded on the surface mesh $\mathcal{M}$. By ensuring that $\mathcal{L}$ satisfies the properties of Definition 3, we can guarantee that we can compute a polycube segmentation from $\mathcal{L}$ at any step. DualCube now proceeds as follows. First, we initialize a loop structure of minimum complexity on $\mathcal{M}$. For a model of genus zero, this is simply a loop structure consisting of three loops (one for each axis label), which corresponds to a single cube as polycube. For models of higher genus, this initialization is more involved (see Section 4.3 for more details). Next, we alter the loop structure by adding or removing loops. In Section 4.1 we explain in detail how to find good loops that improve the quality of the polycube segmentation, and at the same time preserve the characterization of a polycube loop structure stated in Definition 3. Once we have obtained a good loop structure, we compute a corresponding polycube segmentation by placing a corner in every loop region, and connecting the corners of adjacent loop regions via non-intersecting paths on $\mathcal{M}$. We refer to this process as primalization and discuss this step in more detail in Section 4.2. To obtain the best loop structure $\mathcal{L}$, we actually maintain a collection of loop structures, and we utilize a type of evolutionary algorithm to optimize the quality of the corresponding polycube segmentations (see Section 4.3 for details).
Note that DualCube does not need to compute the polycube that corresponds to the polycube segmentation. However, for illustration purposes we frequently show a possible polycube next to the polycube segmentation that our algorithm computes. We construct these polycube examples using a simple heuristic which does not attempt to optimize the exact geometric representation; we omit further details which can, however, be found in our publicly available code, see Section 5.
@@
To find the best loop structure via our evolutionary algorithm, we need to define a measure of quality for the resulting polycube segmentation. There is no universal standard for measuring the quality of a polycube segmentation, as different downstream applications have different requirements. Polycube maps are commonly used in hexahedral meshing, spline fitting, and structured multi-block grid generation, each prioritizing different aspects such as minimal distortion or fewer patches in the segmentation. In general, prior work aims to to balance complexity, often tied to the number of vertices in the corresponding polycube, and distortion, which captures how well the polycube map transforms the original surface into the polycube. Following the terminology introduced by Livesu et al.  [LVS13], we consider a segmentation to be of high quality if it defines a valid polycube topology and achieves a favorable trade-off between fidelity and compactness.
Fidelity
@@
captures the alignment between the segmentation and the input geometry. We measure this via the angular deviation between the normal of each triangle and the axis associated with its assigned label, consistent with the measure used by Dumery et al. [DPM22].
Compactness
reflects the structural simplicity of the segmentation. We measure the compactness of a polycube segmentation via the number of loops in its loop structure.
While Livesu et al. [LVS13] also identify path monotonicity as a desirable property, we do not explicitly optimize for it. Our primalization step results in favorable paths by design, making the addition of a third objective term unnecessary. We opt to retain the simplicity of a two-term formulation focused on fidelity and compactness. In Section 4.3 we discuss which exact metrics we use and how we handle the resulting multi-objective optimization problem.